OpenStudy (anonymous):

Simplify (-12x^4)/(x^4+8x5)

OpenStudy (anonymous):

\[-\frac{ 12 }{ 8x+1 }\]

OpenStudy (anonymous):

what is x not = to

OpenStudy (anonymous):

Distribute the \[-12x ^{4}\] to each of the parts of the other area. So you'll get this as a result \[(x ^{4}*-12x ^{4}) + (8x ^{5}*-12x ^{4})\] (I'm assuming your 8x5 meant an exponent. Then simplify. Do you know how to do that part?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Woah, sorry, ignore that, I just saw that / in there. My fault.

OpenStudy (anonymous):

and yes it did sorry

OpenStudy (anonymous):

and can you tell me is x not = to 0

OpenStudy (anonymous):

or can it be 0

OpenStudy (anonymous):

the question is not done unless it have an equal sign

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is that the right problem

OpenStudy (anonymous):

What you would do is separate the denominator to 2 fractions essentially. So you would have \[\frac{ -12x ^{4} }{ x ^{4} } + \frac{ -12x ^{4} }{ 8x ^{5} }\] You would then simplify. And Marv93w is right, you can't find out what x is equal to unless you have the rest of the equation. It's impossible to solve this, only to simplify it.

OpenStudy (anonymous):

Simplify (-12x^4)/(x^4+8x^5)

OpenStudy (anonymous):

I still need help

OpenStudy (anonymous):

I think I get it. So what your question is is to simplify it and then determine if x can be 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got you started on the simplification, do you know how to do the next step? What do you get if you try it?

OpenStudy (anonymous):

what are the answer choice

OpenStudy (anonymous):

First, do the simplification so you know which set of questions you have. Then, plug in the values given. If you can get a number for a solution then it can be that number. For example, in c, x can't equal 0 because then the denominator would be 0, and you can't divide by 0.

OpenStudy (anonymous):

the answer look like it d if you plug in 1 with the question

OpenStudy (radar):

Marv93w provided you the correct simplification on the first post.

OpenStudy (anonymous):

it is thank you are you sure and @PsychoTink is he right I want for you to tell me is that right is it D

OpenStudy (radar):

1 + 8x must not = 0 So what must x not equal???

OpenStudy (anonymous):

i see it look like it asking for the domian

OpenStudy (anonymous):

@Marv93w how are you saying the answer is D. That's not possible. @s17 the goal of this site (per the code of conduct) is NOT for people to just give the answer, but to help you find the answer on your own and know how to find it.

OpenStudy (radar):

It is not D

OpenStudy (anonymous):

If the denominator is 8x+1, the answer can't be a denominator of 9.

OpenStudy (anonymous):

x=-1/8

OpenStudy (anonymous):

@radar I like your dog and I know I do want help

OpenStudy (anonymous):

this is hard for me I am really good at biology not math

OpenStudy (radar):

Thanks, Marv93w shows you that x can not = - 1/8, because if it does then the denominator becomes 0 and division by 0 is a no no

OpenStudy (anonymous):

okay so is it C

OpenStudy (anonymous):

so you are saying its B

OpenStudy (radar):

Not C

OpenStudy (anonymous):

because if it can not be 1/8 and not 0 its B

OpenStudy (radar):

It is what Marv93w posted in the very first post below your submitted problem.

OpenStudy (anonymous):

you have to put the full question in order to answer the problem

OpenStudy (anonymous):

thanks and again I like your dog I love dogs :) haha thanks again

OpenStudy (anonymous):

Simplify x=7/x^2+4x-21

OpenStudy (radar):